Find the magnetic field 6.15 cm from a long, straight wire that carries a current of 7.84 A .
Answer #1

Magnetic field due to a current carrying wire is given by

B = oI/(2R)

Given, I = 7.84 A; R = 6.15 cm = 0.0615 m

putting all the values in this

B = 4*10-7*7.84/(2*3.14*0.0615)

B = 8.12*10-6 T